Please show me how to solve this problem, I dont know the steps (btw I'm doing quadratic equations)
(3x-3)(4x+2)=0

Answer by richard1234(7193)

Either 3x-3 = 0 or 4x+2 = 0, so x = 1, x = -1/2.